# 2Departament de Física Quàntica i Astrofísica and Institut de Ciències del Cosmos, Universitat de Barcelona, 08028 Barcelona, Spain

### v1

@@ 4.3 Absorption of radiation with large frequency $\omega\gg E_{i}$
As seen in Figure 1, for both the $\ell=1$ photon and the $\ell=2$ graviton, increasing $E_{i}/E_{b}$—that is, approaching the semiclassical black hole regime—leads to a larger absorption cross-section. This is in sharp contrast with the behavior of the $\ell=0$ scalar studied in Emparan:2025sao , where the cross-section decreases as $E_{i}/E_{b}$ increases. The underlying reason is that, for $\ell\geq 1$, the semiclassical absorption cross-section dominates over the quantum one, whereas the opposite holds for $\ell=0$. To illustrate this difference, Figure 4 compares the quantum and semiclassical results for the $\ell=1$ photon. [^9] When $\omega\gg E_{i}$ the quantum result approaches the semiclassical value, since $E_{f}\approx\omega$ is dominated by the strong backreaction of the incoming wave, blurring the distinction between the quantum initial state and its semiclassical limit.
Figure 4: Quantum (4.2) (red) vs. semiclassical (A.3) (black) absorption cross-section of the $\ell=1$ photon as a function of $\omega/E_{i}$, for $E_{i}=0.8\,E_{b}$. The quantum transparency window (4.5) ends at $\omega/E_{i}=0.25$ and is not easily distinguishable. Note how the quantum cross-section is always suppressed compared to the semiclassical one, giving rise to quantum translucency.
The salient feature is that above the transparency window, the quantum cross-section becomes nonzero but remains below the semiclassical one—the black hole is neither transparent nor opaque, but translucent. As we will see, this is due to the still limited density of available final states.
@@
We compare the quantum expressions for the transition probabilities and density of states with their “semiclassical” counterparts (i.e., (C.1) and (C.2) vs. (C.3) and (C.4)). These are labeled “semiclassical” because they result from expanding the quantum formulas in the regime $\omega,E_{b}\ll E_{i}$, where semiclassical physics is expected to hold. However, we use quotation marks advisedly, as their product fails to reproduce the full semiclassical cross-section at frequencies $\omega\gtrsim E_{i}$. For example, in our case study, namely the $\ell=1$ photon with $E_{i}=0.8\,E_{b}$, the quantum cross-section overtakes the product of the “semiclassical” transition probability and density of states once $\omega/E_{i}\gtrsim 2.6$. Therefore the comparison is meaningful only in the regime $\omega/E_{i}\ll 1$.
Figure 5: Quantum vs. semiclassical absorption transition probabilities (left) and density of final states (right) of the $\ell=1$ photon as a function of $\omega/E_{i}$, for $E_{i}=0.8\,E_{b}$ (eqs. (C.1)–(C.4)). The semiclassical density of final states (C.4) has been multiplied by $10^{-2}$ to bring it inside the right panel plot. The transparency window (4.5) ends at $\omega/E_{i}=0.25$. Like for the $\ell=0$ scalar Emparan:2025sao , quantum fluctuations significantly enhance the absorption transitions and strongly suppress the density of final states. These two competing effects determine whether the absorption cross-section is enhanced or suppressed—with the latter being the case for spinning radiation.
Figure 5 illustrates this comparison in the quantum regime $E_{i}\lesssim E_{b}$. It shows the origin of the semiclassical dominance for $\omega\ll E_{i}$: although quantum states absorb spinning radiation more efficiently (left panel), their scarcity relative to semiclassical expectations (right panel) leads to an overall smaller quantum cross-section—unlike the case of scalar s-waves studied in Emparan:2025sao ; Biggs:2025nzs , where quantum effects enhanced absorption. Quantum absorption remains suppressed throughout the full range (4.10) where our formulas apply, creating the phenomenon of quantum translucency.
@@ Appendix B Explicit absorption cross-sections
$f_{\Delta}(E_{i}\pm\omega)$ =±$=\pm\,\frac{\pi^{2}\,\mathcal{B}^{\Delta}_{E_{i}\pm\omega}\,e^{-S_{0}}\,E_{b}^{2\Delta-1}\,\omega}{2^{2\Delta-3}\,\Gamma(2\Delta)\,\left[\cosh\!\left(2\pi\,\sqrt{2\,(E_{i}\pm\omega)/E_{b}}\right)-\cosh\!\left(2\pi\,\sqrt{2\,E_{i}/E_{b}}\right)\right]}\,,$ (C.1)
@@
with $\mathcal{B}^{\Delta}_{E_{i}\pm\omega}$ given by (3.11)–(3.14). If we expand them for $\omega,E_{b}\ll E_{i}$, we find
[…]
$\rho_{\ell}(E_{i}\pm\omega)$ $\approx\frac{(2\ell+1)\,e^{S_{0}}\,e^{4\pi^{2}T/E_{b}}\,e^{\pm\omega/T}}{4\,\pi^{2}\,E_{b}}\,,$ (C.4)
where $T=\beta^{-1}$ is the Hawking temperature of the black hole (3.19).

### v3

@@ 4.3 Absorption of radiation with large frequency $\omega\gg E_{i}$
As seen in Figure 1, for both the $\ell=1$ photon and the $\ell=2$ graviton, increasing $E_{i}/E_{b}$—that is, approaching the semiclassical black hole regime—leads to a larger absorption cross-section. This is in sharp contrast with the behavior of the $\ell=0$ scalar studied in Emparan:2025sao , where the cross-section decreases as $E_{i}/E_{b}$ increases. The underlying reason is that, for $\ell\geq 1$, the semiclassical absorption cross-section dominates over the quantum one, whereas the opposite holds for $\ell=0$. To illustrate this difference, Figure 4 compares the quantum and semiclassical results for the $\ell=1$ photon. [^12] When $\omega\gg E_{i}$ (which is not the semiclassical limit) the quantum result approaches the semiclassical value, since $E_{f}\approx\omega$ is dominated by the strong backreaction of the incoming wave, blurring the differences between the quantum initial state and its semiclassical limit.
Figure 4: Quantum (4.2) (red) vs. semiclassical (A.3) (black) absorption cross-section of the $\ell=1$ photon as a function of $\omega$, for $E_{i}=0.8\,E_{b}$. The quantum transparency window (4.5) ends at $\omega/E_{i}=0.25$ and is not easily distinguishable. Note how the quantum cross-section is always suppressed compared to the semiclassical one, giving rise to quantum translucency.
The salient feature is that above the transparency window, the quantum cross-section becomes nonzero but remains below the semiclassical one—the black hole is neither transparent nor opaque, but translucent. As we will see, this is due to the still low density of available final states.
@@
In Figure 5 we compare the quantum expressions for the transition probabilities and density of states with their semiclassical counterparts (i.e., (C.1) and (C.2) vs. (C.3) and (C.4)), in the quantum regime $E_{i}\lesssim E_{b}$ (where, recall, there is no stimulated emission of spin).
Figure 5: Quantum vs. semiclassical absorption transition probabilities (left) and density of final states (right) of the $\ell=1$ photon as a function of $\omega$, for $E_{i}=0.8\,E_{b}$ (eqs. (C.1)–(C.4)). The transparency window (4.5) ends at $\omega/E_{i}=0.25$. As in the case of the $\ell=0$ scalar Emparan:2025sao , quantum fluctuations markedly enhance absorption transitions while sharply suppressing the density of final states. These competing effects determine whether the absorption cross-section is enhanced or suppressed, the latter being the case for spinning radiation. We plot $f_{3}$ in units of $e^{-S_{0}}E_{b}^{6}$, and $\rho_{1}$ in units of $e^{S_{0}}E_{b}^{-1}$. The log-scale vertical axes highlight the large disparity between semiclassical and quantum values.
This comparison reveals the origin of the semiclassical dominance: while quantum states absorb spinning radiation more efficiently (left panel), their scarcity relative to semiclassical expectations (right panel) results in a lower overall quantum cross-section—in contrast to the case of scalar s-waves studied in Emparan:2025sao ; Biggs:2025nzs , where quantum effects actually enhanced absorption. In the present case, quantum absorption remains suppressed throughout the entire range (4.10) where our expressions are valid, creating the phenomenon of quantum translucency.
@@ Appendix B Explicit absorption cross-sections
$f_{\Delta}(E_{i}\pm\omega)$ $=\pm\,\frac{\pi^{2}\,\mathcal{B}^{\Delta}_{E_{i}\pm\omega}\,e^{-S_{0}}\,E_{b}^{2\Delta-1}\,\omega}{2^{2\Delta-3}\,\Gamma(2\Delta)\,\left[\cosh\!\left(2\pi\,\sqrt{2\,(E_{i}\pm\omega)/E_{b}}\right)-\cosh\!\left(2\pi\,\sqrt{2\,E_{i}/E_{b}}\right)\right]}\,$, (C.1)
@@
with $\mathcal{B}^{\Delta}_{E_{i}\pm\omega}$ given by (3.11)–(3.15). If we expand them in the semiclassical regime $\omega,E_{b}\ll E_{i}$, we find
$f_{\Delta}(E_{i}\pm\omega)$ $\approx\pm\frac{2\,(2\pi)^{2\Delta}\,e^{-S_{0}}\,E_{b}\,\sinh\!\left(\!\frac{\beta\omega}{2}\!\right)\,\Gamma\!\left(\!\Delta+i\,\frac{\beta\omega}{2\pi}\!\right)\,\Gamma\!\left(\!\Delta-i\,\frac{\beta\omega}{2\pi}\!\right)}{\Gamma(2\Delta)\,\beta^{2\Delta-1}\,e^{\frac{(2\pi)^{2}}{\beta E_{b}}}\,\left(e^{\pm\beta\omega}-1\right)}\,$, (C.3)
$\rho_{\ell}(E_{i}\pm\omega)$ $\approx\frac{(2\ell+1)\,e^{S_{0}}}{(2\pi)^{2}\,E_{b}}\,e^{\frac{(2\pi)^{2}}{\beta E_{b}}\pm\beta\omega}\,$, (C.4)
where $\beta$ is the inverse Hawking temperature of the black hole (3.21). In writing the expression for the semiclassical density of states (C.4), we also assumed $\omega\gg E_{b}$ since semiclassical physics is insensitive to the quantum energy gap $E_{0,\ell}\sim E_{b}$. However, note that we do not need to assume $\omega\gg E_{b}$ in order to obtain the semiclassical greybody factor (3.24). If we do not make this assumption, then we would have an extra exponential $e^{-\beta E_{0,\ell}}$ in the density of states (C.4).